# HMT: Hierarchical Memory Transformer for Long Context Language Processing

### v2

@@ Appendix D Ablation Study of HMT Memory Recall Behavior
One insight of using memory recall in HMT is handling frequent context switching to previously discussed topics or new topics. To evaluate this property, we employ PubMedQA and artifact the dataset with multiple contexts, mentioned in Section 6.1. In this section, we will discuss other dataset manipulations on PG-19 to investigate the memory recall behavior of HMT further.
@@
One way to manually introduce context switching is by interleaving the samples. For every 2 samples in the PG-19 dataset, we alternatively concatenate a segment of 256 tokens in each sample together to create a new sample. Therefore, a context switch will be invoked every 256 tokens. We fine-tuned and benchmarked HMT with Llama 2 7B over the artifact dataset. As a result, HMT can enhance the effectiveness of the baseline Llama 2 model, while RMT will worsen it, as shown in Figure 14. We record the context distance of memory recall for 30k-token input, illustrated in Figure 14, and notice a periodical recall distribution, indicating that HMT can capture the context-switching pattern.
@@
Figure 13: Effectiveness of HMT and RMT with Llama 2 7B evaluated over PG-19 with interleaving sample. HMT is 12.02% better than RMT in terms of PPL for 2k to 100k-token samples.
@@
Figure 14: Histogram of context distance between the current segment and the segment corresponding to the memory embedding with the highest softmax score in the memory recall mechanism. The dataset evaluated is the PG-19 with interleaving samples.
@@
To demonstrate that HMT’s behavior is aligned with the context-switching pattern, we further manipulate the PG-19 dataset by inserting 256 “\$” tokens for every 256 tokens to dilate each sample. Intuitively, the segment containing “\$” should be considered as irrelevant information and recalled infrequently. Figure 15 shows the memory recall pattern of HMT with Llama 2 7B over the dilated PG-19 dataset. We observe that HMT not only exhibits a periodical recall pattern but also successfully captures the position of irrelevant segments and avoids recalling them.
@@
Figure 15: Histogram of context distance between the current segment and the segment corresponding to the memory embedding with the highest softmax score in the memory recall mechanism. The dataset evaluated is the dilated PG-19 dataset. Each sample is 25.6k tokens.
Appendix E Relationships Between Effectiveness and Size of Sensory Memory
@@
Table 7: Effectiveness of HMT + Llama 2 7B evaluated on Wikitext-103 with 100k-token samples, with a various sensory memory size. The segment size is 256 tokens. The effectiveness improves and then degrades with an increasing number of embeddings that are preserved for sensory memory.
@@
Appendix F Dataset Split for PubMedQA
@@
Appendix G Hardware Acceleration Consideration
The design of HMT can be implemented efficiently with customizable hardware, such as Field-Programmable Gate Array (FPGA). In contrast to GPUs, FPGA has two advantages in constructing accelerator HMT with high speed and low energy consumption:
FPGA has a flexible architecture. Users can define the FPGA fabric to have data-driven modules for stream architecture. Processing elements (PE) consume data from streams asynchronously and initiate computation once required data are ready. Stream architecture is naturally favored by the recurrent sequence models. Previous works in building an RNN accelerator (Chang & Culurciello, 2017; Khoda et al., 2023) and stream overlay for LSTM (Ioannou & Fahmy, 2022) on FPGA demonstrate a 1.25x to 16x throughput advantage over GPUs for low-batch tasks. Furthermore, the reconfigurable architecture allows fabric designers to save energy by only allocating necessary resources, instead of enforcing warp-level execution in GPUs (Abdelkhalik et al., 2022).
FPGA has sophisticated support in task-level parallelism. Frameworks such as TAPA (Guo et al., 2023) and PASTA (Khatti et al., 2023) generate multi-module FPGA designs with a single source file such that each function runs independently and communicates with FIFOs. On the contrary, the most popular GPU acceleration toolkit, CUDA, only has abstractions for task-level pipelining using stream (Li et al., 2014).
Based on the advantages of FPGA, we proposed the following optimizations for HMT:
Intramodel Pipelining for Batched Inputs. The transformer-based backbone model is naturally layered by multiple transformer blocks and the computation of each layer only depends on the previous layer. Therefore, the stream of inputs can be processed in the pipeline. Moreover, since the HMT processes long inputs in segments, the computation can further be pipelined with batches of inputs by round-robin scheduling.
Task-Level Parallelism. Although recalling memory requires the memory token from the inference of the previous segment, overlapping both tasks with an extra low-cost cross-attention computation is possible. When the model performs a forward pass on the current segment, the representation extraction of the next segment can generate the embedding $H_{n}^{sum}$ and initiate the cross-attention with previous memory tokens for segments except the current one. The model will then wait until the memory token $H^{mem}_{n}$ of the current segment is produced and compute the attention score between $H_{n+1}^{sum}$ and $H^{mem}_{n}$. Finally, the attention scores will be used to calculate the output embedding as the memory token for the next segment. Specifically, let $H_{[n-N+1,n)}^{mem}$ be the embedding of memory tokens except for the current segment, $H_{n}^{mem}$ be the memory token embedding of the current segment, and $H_{n+1}^{sum}$ be the summary embedding of the next segment, the computation
$K=H_{[n-N+1,n)}^{mem}W_{k},Q=H_{n+1}^{sum}W_{q},A=QK^{T}$
can be overlapped with the forward pass of the previous segment and the computation
$K_{n}=H_{n}^{mem}W_{k},A_{n}=QK_{n}^{T}$
$H^{S}_{n}=\text{Softmax}(\frac{[A\circ A_{n}]}{\sqrt{d_{h}}})[H_{[n-N+1,n)}^{mem}\circ H_{n}^{mem}]$
will run in serial. Table 8 illustrates the runtime breakdown of each computation phase, proving that the latency of representation abstraction and computing cross-attention without the last memory token can be fully covered by the forward pass of the previous segment.
We leave the actual hardware implementation on an FPGA to future work.
Table 8: GPU Runtime breakdown of each step in memory recall mechanism of HMT per segment of inputs. The segment length is 256 and the backbone model is OPT 2.7B.
Step Runtime (ms)
Forward pass of the segment 85.18
Representation extraction 49.20
Compute $A$ 1.38
Compute $A_{n}$ 0.25
Appendix H Gradient Stability in HMT and RMT
@@
Figure 16: Backward propagation flows of HMT and RMT. The gradient of the first memorization prompt embedding (the red block on the right of the first segment) has multiple branches through the memory recall unit. On the contrary, the gradient for RMT requires propagating through each segment.
@@
In this section, we will formulate a mathematical description for the gradient stability of HMT when training with BPTT. BPTT with RMT behaves similarly to RNN which suffers from a vanishing or exploding gradient when there is a long chain of the gradient graph after unrolling (Pascanu et al., 2013). Specifically, for a generic RNN model with the following form:
@@
Whether the gradient will explode or vanish depends on the input distribution and the function $F_{t}$. If $\forall x_{t},\frac{\partial F}{\partial m_{t}}(x_{t})>0$, then the gradient will explode. Otherwise if $\forall x_{t},\frac{\partial F}{\partial m_{t}}(x_{t})<0$, the gradient vanishes. Consequently, training RMT with a very high BPTT unroll depth can be inefficient. For HMT, with the assistance of the memory recall mechanism, the gradient is not prone to explosion or vanishing. Intuitively, the backward propagation of HMT for the memorization prompt embedding contains multiple short sub-branches to prevent gradient vanishing and the memory recall mechanism can modulate the propagation chain to avoid gradient explosion, depicted in Figure 16. Let $G_{t}(z_{t},m_{t},m_{t-1},\dots,m_{1})=m^{\prime}_{t}$ be the memory search function where $z_{t}$ is the representation extraction of segment at time $t$. Let $s$ be the summarization token for representation extraction. The gradient for HMT is
@@
Appendix I Distributed Training with Memory Consumption Optimization
Although Bulatov et al. proves that unrolling more segments can improve the model effectiveness, they limit the number of segments unrolled to 4 with 2 NVIDIA A100 80GB GPUs since the maximum BPTT unroll depth is bounded by the GPU VRAM limit. There are three sources of VRAM consumption: model parameters, intermediate data (input segments, long-term memory, raw outputs of each segment, etc.), and optimization data (gradient and optimizer states). Although the computations of later segments do not require the intermediate data from the previous segment, the original BPTT will keep them on GPU by default. To reduce memory consumption, we customize the program to offload and load intermediate data for each input segment between the CPU and GPUs and distribute optimizer states and gradients throughout multiple GPUs running Zero Redundancy Optimizer (ZeRO) (Rajbhandari et al., 2020) Stage 2 in DeepSpeed (Rasley et al., 2020). These allow the model to unroll up to 15 segments with HMT. To train larger models, we employ LoRA (Hu et al., 2021) with rank 8. This allows us to fit 7B models to 4 MI210 GPUs.

### v3

@@ Appendix G HMT Memory Retrieval Behavior
One insight of using memory retrieval in HMT is handling frequent context switching to previously discussed topics or new topics. To evaluate this property, we employ PubMedQA and artifact the dataset with multiple contexts, mentioned in Section 5.1. In this section, we will discuss other dataset manipulations on PG-19 to investigate the memory retrieval behavior of HMT further.
@@
One way to manually introduce context switching is by interleaving the samples. For every 2 samples in the PG-19 dataset, we alternatively concatenate a segment of 256 tokens in each sample together to create a new sample. Therefore, a context switch will be invoked every 256 tokens. We fine-tuned and benchmarked HMT with Llama 2 7B over the artifact dataset. As a result, HMT can enhance the effectiveness of the baseline Llama 2 model, while RMT will worsen it, as shown in Figure 14. We record the context distance of memory retrieval for 30k-token input, illustrated in Figure 14, and notice a periodical recall distribution, indicating that HMT can capture the context-switching pattern.
@@
Figure 13: Effectiveness of HMT and RMT with Llama 2 7B evaluated over PG-19 with interleaving samples. HMT is 12.02% better than RMT in terms of PPL for 2k to 100k-token samples.
@@
Figure 14: Histogram of context distance between the current segment and the segment corresponding to the memory embedding with the highest softmax score in the memory retrieval mechanism. The dataset evaluated is the PG-19 with interleaving samples.
@@
To demonstrate that HMT’s behavior is aligned with the context-switching pattern, we further manipulate the PG-19 dataset by inserting 256 “\$" tokens for every 256 tokens to dilate each sample. Intuitively, the segment containing “\$" should be considered as irrelevant information and recalled infrequently. Figure 15 shows the memory retrieval pattern of HMT with Llama 2 7B over the dilated PG-19 dataset. We observe that HMT not only exhibits a periodical recall pattern but also successfully captures the position of irrelevant segments and avoids recalling them.
@@
Figure 15: Histogram of context distance between the current segment and the segment corresponding to the memory embedding with the highest softmax score in the memory retrieval mechanism. The dataset evaluated is the dilated PG-19 dataset. Each sample is 25.6k tokens.
Appendix H Relationships Between Effectiveness and Size of Sensory Memory
@@
Table 8: Effectiveness of HMT + Llama 2 7B evaluated on Wikitext-103 with 100k-token samples, with various sensory memory sizes. The segment size is 256 tokens. The effectiveness improves and then degrades with an increasing number of embeddings that are preserved for sensory memory.
@@
Appendix I Dataset Construction for PubMedQA
@@
We artifact the long-context dataset from PubMedQA as the following: (1) select $M$ question-context-answer tuples from the dataset. Let this set of tuples be $\{(C_{0},Q_{0},A_{0}),(C_{1},Q_{1},A_{1}),\dots,(C_{T},Q_{T},A_{T})\}$, where $C_{n}$ are contexts, $Q_{n}$ are questions, $A_{n}$ are answers for $0\geq n\geq M$. Answers can be either long answers with detailed reasoning or short answers (either “yes", “no", or “maybe"). (2) Concatenate all the contexts from each tuple to form a long context and append the questions and answers for each tuple. This will create $M$ sequences: ${C_{0}C_{1}\dots C_{T}Q_{0}A_{0}}$, ${C_{0}C_{1}\dots C_{T}Q_{1}A_{1}},\dots$, ${C_{0}C_{1}\dots C_{T}Q_{T}A_{T}}$. By controlling the value of $M$, we can determine the fraction of useful information in the context for each question and better understand the filtering and selection ability of HMT and the baseline model.
Appendix J Gradient Stability in HMT and RMT
Both HMT and RMT are trained using backward propagation through time (BPTT) Mozer (2013), a technique utilized to train the RNN model by unrolling recurrent forward passes of the model to optimize long-sequence learning. One issue with RMT training with BPTT is the gradient explosion and vanishing problem. With a higher BPTT unroll depth, the effectiveness of RMT will first increase and then decrease, with a slow reduction or even an increase in training loss. As seen in Figure 16, we use the Wikitext-103 dataset with various BPTT unroll depths to access the effectiveness of RMT with the OPT 2.7B backbone model. For both 2k and 10k token inputs, we observe a rising PPL when unrolling more than 5 segments during training.
Figure 16: Effectiveness of training RMT with BPTT with different unroll depths for 2K tokens and 10K tokens input from the Wikitext-103 dataset. The backbone model is OPT 2.7B, with 256 tokens per segment during inference.
Table 9: Relationship between the BPTT unroll depth and the test PPL of Wikitext-103 for OPT 2.7B with HMT. The experiment is evaluated on samples with 10k tokens. HMT preserved 32 tokens from the previous segment as the sensory memory and saved 300 memory embeddings for the memory retrieval. The segment size is 256 tokens.
BPTT Unroll Depth Test PPL (Wikitext)
2 9.36
5 9.15
15 8.20
Unlike RMT, HMT does not suffer from gradient vanishing or explosion as BPTT unroll depth increases due to the memory retrieval mechanism. Table 9 reveals that HMT can improve its effectiveness continuously as the BPTT unroll depth increases during training. Therefore, HMT will be more effective when the BPTT unroll depth increases. A detailed gradient stability analysis is presented in Appendix J. Furthermore, we applied several techniques to optimize the GPU memory consumption to increase the maximum trainable BPTT unroll depth compared with RMT, described in Appendix K.
@@
Figure 17: Backward propagation flows of HMT and RMT. The gradient of the first memorization prompt embedding (the red block on the right of the first segment) has multiple branches through the memory retrieval unit. Where the HMT gradient does not require propagation through each segment, the RMT gradient does.
@@
In this section, we will formulate a mathematical description for the gradient stability of HMT when training with BPTT. BPTT with RMT behaves similarly to RNN which suffers from a vanishing or exploding gradient when there is a long chain of the gradient graph after unrolling Pascanu et al. (2013). Specifically, for a generic RNN model with the following form:
@@
Whether the gradient will explode or vanish depends on the input distribution and the function $F_{t}$. If $\forall x_{t},\frac{\partial F}{\partial m_{t}}(x_{t})>0$, then the gradient will explode. Otherwise if $\forall x_{t},\frac{\partial F}{\partial m_{t}}(x_{t})<0$, the gradient vanishes. Consequently, training RMT with a very high BPTT unroll depth can be inefficient. For HMT, with the assistance of the memory retrieval mechanism, the gradient is not prone to explosion or vanishing. Intuitively, the backward propagation of HMT for the memorization prompt embedding contains multiple short sub-branches to prevent gradient vanishing and the memory retrieval mechanism can modulate the propagation chain to avoid gradient explosion (Figure 17). Let $G_{t}(z_{t},m_{t},m_{t-1},\dots,m_{1})=m^{\prime}_{t}$ be the memory search function where $z_{t}$ is the representation extraction of segment at time $t$. Let $s$ be the summarization token for representation extraction. The gradient for HMT is
@@
Appendix K Distributed Training with Memory Consumption Optimization
Although Bulatov et al. (2022) proves that unrolling more segments can improve the model effectiveness, they limit the number of segments unrolled to 4 with 2 NVIDIA A100 80GB GPUs since the maximum BPTT unroll depth is bounded by the GPU VRAM limit. There are three sources of VRAM consumption: model parameters, intermediate data (input segments, long-term memory, raw outputs of each segment, etc.), and optimization data (gradient and optimizer states). Although the computations of later segments do not require the intermediate data from the previous segment, the original BPTT will keep them on GPU by default. To reduce memory consumption, we customize the program to offload and load intermediate data for each input segment between the CPU and GPUs and distribute optimizer states and gradients throughout multiple GPUs running Zero Redundancy Optimizer (ZeRO) Rajbhandari et al. (2020) Stage 2 in DeepSpeed Rasley et al. (2020). These allow the model to unroll up to 15 segments with HMT. To train larger models, we employ LoRA Hu et al. (2021) with rank 8. This allows us to fit 7B models to 4 MI210 GPUs.
Appendix L License and Links of Datasets and Models
Datasets:
Models:
All datasets and models are publicly available and free to use.